solve the following proportion for y
11/ 7 = y /5

Answer by checkley77(12844)

11/ 7 = y /5 CROSS MULTIPLY
7Y=5*11
7Y=55
Y=55/7 OR 7.857 ANS.
PROOF:
11/7=(55/7)/5
11/7=55/7*1/5
11/7=55/35
11/7=11/7